How many liters of a 14% alcohol solution must be mixed with 20 liters of a 50% solution to get a 30% solution?
Answer by checkley75(3666)

.14x+.50*20=.30(x+20)
.14x+10=.3x+6
.14x-.3x=6-10
-.16x=-4
x=-4/-.16
x=25 liters of 14% alcohol is needed to make the required mix
proof
.14*25+.5*20=.3(25+20)
3.5+10=.3(45)
13.5=13.5